OpenStudy (anonymous):

A rectangular field is enclosed on 3 sides by a fence, while the 4th side is a stream. The fence opposite the stream is 150 m longer than the other sides. If the area of the field is 70,000 m2, find its dimensions.

OpenStudy (campbell_st):

 let the width of the field be x then the length is x + 150 Area = Length x width so \[70000 = x(x + 150) .... or.... x^2 + 150x - 70000 = 0\] solve the quadratic to find the dimensions. factorising (x + 350)(x - 200) = 0 then x = -350 or 200 only consider the positive measurement so the width is 200 metres and the length is 350 metres